a+b=2 than a=4 , b=?

Grade:12

10 Answers

venky
8 years ago
b=2
RAJU
8 years ago
b=-2
santhosh
8 years ago
b=-2
Nihal Patel
8 years ago
a + b = 2
a = 4 
b = ?
Substitute the value of a in the equation:
4 + b = 2 
b = 2 – 4
Therefore, b = -2
SEELA YUVARAJU
8 years ago
-2
R.GANGARAJU
8 years ago
-2
ANJI BABU
8 years ago
-2
Raghu Vamshi Hemadri
8 years ago
 
a + b = 2
a = 4 
b = ?
Substitute the value of a in the equation:
4 + b = 2 
b = 2 – 4
Therefore, b = -2
Prabhakar ch
8 years ago
-2
AYESHA
8 years ago
Hi...................................b=-2